OpenStudy (anonymous):

dianne pays $28 to enter the state fair plus $2 for each ride what equation woul you use to repersent her cost

OpenStudy (anonymous):

A.y=28x,B.y=2x+23,C.y=28x+2,D.y=2x+30

OpenStudy (anonymous):

Cost=28+2x or B.

OpenStudy (anonymous):

you would use 28+2x because you only pay $28 once, and for each ride, which is x, you pay $2 soo...

OpenStudy (anonymous):

ITS B thanks